OpenStudy (anonymous):

Find all the zeros of the equation. -3x^4+27x^2+1200=0

OpenStudy (mathstudent55):

Try a substitution. Let \(u = x^2\). Then, \(u^2 = x^4\). Make those substitutions. Then you'll have a quadratic equation in u. Solve for u. Also notice that -3, 27, and 1200 are all divisible by -3.